in a 10-kilometer race, there are 112 more men entered than women. Find the number of female runners if the total number of runners in the race is 600
Answer by stanbon(75887)

in a 10-kilometer race, there are 112 more men entered than women. Find the number of female runners if the total number of runners in the race is 600
--------
m = w + 112
-----
m + w = 600
-------
Substitute for "m" and solve for "w":
w+112 + w = 600
2w = 488
w = 244 (# of women)
----
m = w + 112
m = 244+112 = 356 (# of men)
==============================
Cheers,
Stan H.
==============================